OpenStudy (anonymous):

Static equilibrium; A weight hangs from a long board of negligible weight. Find the magnitude of the tension T in the horizontal rope and the horizontal and vertical components of force exerted on the board at P.

OpenStudy (anonymous):

 \[\sum \tau = 0\] \[\sum F _{y} = 0\] \[\sum F _{x} = 0\]